# Marginally interpretable spatial logistic regression with bridge processes

### v1

@@ 5 Childhood malaria prevalence in the Gambia
We illustrate the proposed model by analyzing data on malaria among children in Gambia (Thomson et al.,, 1999), The dataset, available in R package geoR (Diggle et al.,, 1998), contains $\sum_{i=1}^{n}N_{i}=2035$ children’s malaria infection statuses from $n=65$ villages, along with covariates including age, bed net use, net treatment with insecticides, a satellite-derived measure of greenness, and the presence of a health center in the villages. The study of population-averaged associations between malaria prevalence and covariates is useful for nationwide decision-making; for example, on supplying bed nets. Prediction of malaria prevalence at new locations is also of substantial interest. Previous literature suggests residual spatial dependence in these data (Diggle et al.,, 2002; Bai et al.,, 2014; Cattelan and Varin,, 2018).
@@
We choose the same set of variables as Cattelan and Varin, (2018), reproducing their results based on spatial generalized estimating equations, and comparing with bridge and Gaussian process random effect logistic models with an exponential correlation kernel. We use the same priors as in the simulation study (Gelman,, 2006; Gelman et al.,, 2008). The Markov chains using Algorithm 1 show excellent convergence with running time of around 2 mins for a total of 11,000 iterations under Apple M1 processor. See Appendix A.5 for detailed settings.
@@
The results are summarized in Table 2. The site-specific estimates using the bridge and Gaussian process random effect models give almost identical results. From the bridge process model, the posterior mean of the attenuation parameter is $\hat{\phi}=0.75$ with 95% credible interval $(0.46,0.97)$. The spatial generalized estimating equation and the proposed method give similar results for population-averaged estimates, except for the intercept and greenness. Specifically, the significance of the intercept and greenness coefficients at the 0.05 level differs across two methods. This discrepancy can be partly explained by simulation results, where the coverage of the intercept estimate $\hat{\beta}^{M}_{0}$ under the true $\beta_{0}^{M}=0$ is lower than the nominal level with the spatial generalized estimating equation method but not with the proposed method; see Table A.1 in Appendix A.4.
@@ Acknowledgement
This research was partially supported by the National Institutes of Health (grant ID R01ES035625), by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 856506), and by the Office of Naval Research (N00014-21-1-2510). The authors thank Amy Herring and Georgia Papadogeorgou for helpful discussions.
@@
Section A.1 contains proofs of the statements presented in the main article. Section A.2 describes the posterior inference algorithm and associated computational strategies. Section A.3 contains details of the multivariate bridge distribution, including a mixture representation. Sections A.4 and A.5 provide further details of the simulation study and the analysis of the Gambia malaria data.
@@ A.2.4 Details of scalable computation with low-rank dependence structure
When $n$ is large, a low-rank correlation kernel can significantly reduce the computational burden. We choose a kernel based on the modified predictive process (Finley et al.,, 2009),
~$\tilde{\mathcal{R}}(s,s^{\prime})=r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})+1(s=s^{\prime})\{1-r(s)^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s^{\prime})\}$
where $\{\tilde{s}_{1},\dots,\tilde{s}_{q}\}\subset\mathscr{S}$ are knots, $q\ll n$, $r(s)=[\mathcal{R}(s,\tilde{s}_{k})]_{k=1}^{q}$ and $R_{qq}=[\mathcal{R}(\tilde{s}_{k},\tilde{s}_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ is a $q\times q$ matrix from the parent kernel. Under this low-rank correlation kernel, denoting $R_{nq}=[\mathcal{R}(s_{i},\tilde{s}_{k})]_{i=1,k=1}^{n,q}$ as an $n\times q$ matrix and $R_{qn}=R_{nq}^{\mathrm{\scriptscriptstyle T}}$, we have a low-rank structured correlation matrix for $n$ realizations $\tilde{R}=[\tilde{R}(s_{i},s_{i^{\prime}})]_{i,i^{\prime}=1}^{n,n}=R_{nq}R_{qq}^{-1}R_{qn}+D_{nn}$ where $D_{nn}$ is a diagonal matrix with elements $1-r(s_{i})^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s_{i})$ for $i=1,\dots,n$. First, step 1 of Algorithm 1 involves $n\times n$ matrix inversion in equation (A.1). This can be reduced to a $q\times q$ matrix inversion problem using the Woodbury matrix identity,
@@ A.3 Details of multivariate bridge distribution
A.4 Details on simulation study
We first describe details of prior specifications for the bridge process and Gaussian process random effect logistic models. Following Gelman et al., (2008), we choose Cauchy priors for $\beta$, specifically location 0 scale 10 for $\beta_{0}$ and location 0 scale 1.25 for $\beta_{1}$. For the random effect standard deviation $\sigma_{u}$, we choose a weakly informative half-Cauchy prior (Gelman,, 2006), namely $p(\sigma_{u})=2/\{\pi(1+\sigma_{u}^{2})\}$. For the bridge process models where $\sigma_{u}=3^{-1/2}\pi(\phi^{-2}-1)^{1/2}$, this becomes $p(\phi)=(12)^{1/2}/[\{\pi^{2}-(\pi^{2}-3)\phi^{2}\}(1-\phi^{2})^{1/2}]$ for $\phi\in(0,1)$. For the prior of spatial range $\rho$ in Matérn correlation kernel with known smoothness 1.5, we assign uniform prior $\rho\sim\mathrm{Unif}(0.001,0.3)$ for both bridge and Gaussian process random effect models, which corresponds to effective range approximately between 0.00475 and 1.425 in unit square domain.
We next describe detailed settings of the inference procedure. For the spatial generalized estimating equation method of Cattelan and Varin, (2018) with the same notations therein, we used the number of bins $B=13$, radius $h=0.05$, and $d_{\mathrm{max}}=0.3$ to obtain the empirical spatial lorelogram, and optimize the parameters $\alpha_{2}$ (sill) and $\alpha_{3}$ (range) of Matérn kernel with smoothness 1.5 without nugget. For Step 2(i) of the bridge process random effect model, we set the number of particles $L=20$, use coordinatewise logit transform to map parameters $(\phi,\rho)\in(0,1)\times(0.001,0.3)$ to $\mathbb{R}^{2}$, and utilize Metropolis-Hastings proposal adaptation settings as in Haario et al., (2001). For the bridge and Gaussian process random effect models, we run 11,000 iterations, with the first 1,000 samples discarded as burn-in and 1,000 samples are saved with 10 thin-in rates. With $n=200$ spatial locations without low-rank structure and total number of binary data of $N=2000$, the running time of Markov chain Monte Carlo based on a pure R implementation is about 38 mins for the bridge process random effect model and about 20 mins for the Gaussian process random effect model under the Intel(R) Xeon(R) Gold 6336Y 2.40GHz CPU environment.
Table A.1: Comparison of population-average $(\hat{\beta}_{0}^{M})$ and site-specific $(\hat{\beta}_{0})$ estimates of intercept across three methods along with coverage probabilities at nominal level 0.95.
Data from bridge process Data from Gaussian copula process
Population-average Site-specific Population-average Site-specific
$\rho$ Method Sp GEE BrP RE BrP RE GP RE Sp GEE BrP RE BrP RE GP RE
$0.05$ Bias 0.001 -0.003 -0.021 -0.022 -0.005 -0.005 -0.007 -0.005
RMSE 0.237 0.245 0.413 0.412 0.227 0.237 0.339 0.338
Coverage 91.5% 96.0% 96.0% 96.0% 92.0% 96.5% 96.5% 95.5%
$0.1$ Bias -0.013 -0.002 -0.019 -0.019 -0.011 0.003 0.008 0.006
RMSE 0.386 0.382 0.621 0.617 0.380 0.377 0.568 0.568
Coverage 85.5% 97.5% 97.5% 97.0% 85.0% 99.0% 99.0% 98.5%
Sp GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP RE, proposed bridge process random effect model; GP RE, Gaussian process random effect model; RMSE, root mean squared error.
In addition to the results of $\hat{\beta}_{1}^{M}$ and $\hat{\beta}_{1}$, we present a summary of comparisons of $\hat{\beta}_{0}^{M}$, $\hat{\beta}_{0}$ and predictive performance. The intercept coverage probability with true $\beta_{0}^{M}=\beta_{0}=0$ under the spatial generalized estimating equation is slightly lower than the nominal level as the spatial dependence increases, which is consistent with the findings of Cattelan and Varin, (2018). The site-specific estimates of the bridge and Gaussian process random effect models are essentially the same. In terms of predictive performance based on held-out data of size $n_{\mathrm{test}}=50$, $N_{\mathrm{test}}=500$, and test covariates generated as in the training case, we analyze the test log-likelihood averaged over $N_{\mathrm{test}}=500$ binary data and area under the receiver operating characteristic curve. For the spatial generalized estimating equation method, we used $\mathrm{logit}^{-1}(x_{ij}^{{\mathrm{\scriptscriptstyle T}}}\hat{\beta}^{M})$ as predicted probabilities and for Bayesian approaches, we used the posterior predictive mean. The results suggest that the bridge process and Gaussian process random effect models produce virtually identical predictions which are superior to marginal methods that cannot yield site-specific predictions.
Table A.2: Comparison of predictive performance across three methods based on held-out binary data, averaged across 200 replicates.
Data from bridge process Data from Gaussian copula process
$\rho$ Method Sp GEE BrP RE GP RE Sp GEE BrP RE GP RE
$0.05$ Test log-likelihood ×$\times 10^{2}$63.446 -53.073 -53.069 -63.545 -52.564 -52.572
Area under curve ×$\times 10^{2}$ 68.304 80.130 80.137 68.280 80.866 80.861
$0.1$ Test log-likelihood ×$\times 10^{2}$62.390 -49.963 -49.963 -62.561 -49.108 -49.103
Area under curve ×$\times 10^{2}$ 68.541 81.931 81.932 68.520 83.033 83.037
Sp GEE, spatial generalized estimating equation method of Cattelan and Varin, (2018); BrP RE, proposed bridge process random effect model; GP RE, Gaussian process random effect model. The Monte Carlo standard errors are all less than 0.5 for $\rho=0.05$ and all less than 0.7 for $\rho=0.1$.
A.5 Details of the Gambia malaria data case study
For the spatial generalized estimating equation method, we reproduce the result of Cattelan and Varin, (2018) using the same settings based on an exponential spatial kernel. In terms of prior specifications for Bayesian approaches, we standardize the covariates following the suggestion of Gelman et al., (2008) and assign Cauchy priors for $\beta$, specifically location 0 scale 10 for $\beta_{0}$ and location 0 scale 2.5 for non-intercept coefficients. For the random effect standard deviation $\sigma_{u}$, we assign a weakly informative half-Cauchy prior (Gelman,, 2006), as in the simulation studies. For the prior on the spatial range $\rho$ in the Matérn correlation kernel with known smoothness 0.5, namely $\mathcal{R}(s,s^{\prime})=\exp(\|s-s^{\prime}\|_{2}/\rho)$, we assign a uniform prior $\rho\sim\mathrm{Unif}(0.01,100)$ for both the bridge and Gaussian process random effect models. We run three Markov chains with 11,000 iterations, where the first 1,000 samples are discarded as burn-in and 10,000 samples are saved without thinning. The running time of the algorithm based on a pure R implementation is about 2 mins for the bridge process random effect model and about 1.5 mins for the Gaussian process random effect model under the Apple M1 3.20GHz CPU environment.
Trace plots of the bridge process and Gaussian process random effect models are shown in Figure A.2, which show excellent convergence including for the the intercept term. In addition to the estimated population-averaged and site-specific coefficients, we have posterior mean $\hat{\rho}=33.3(14.2,47.2)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the bridge process, posterior mean $\hat{\rho}=29.9(13.4,40.1)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the Gaussian process random effect models, where parenthesis corresponds to the posterior interquartile range. The widely applicable information criterion (Gelman et al.,, 2014), conditional on random effects, since there are no replicates, gives 2329.5 (40.2) for the bridge process random effect model and 2326.3 (40.0) for the Gaussian process random effect model, where parenthesis correspond to standard errors.
Code to reproduce the above analyses is available at https://github.com/changwoo-lee/spbridge.

### v2

@@ 5 Childhood malaria prevalence in Gambia
We illustrate the proposed model by analyzing malaria data among children in Gambia (Thomson et al.,, 1999), The dataset, available in R package geoR (Diggle et al.,, 1998), contains $\sum_{i=1}^{n}N_{i}=2035$ children’s malaria infection statuses from $n=65$ villages, along with covariates including age, bed net use, net treatment with insecticides, a satellite-derived measure of greenness, and the presence of a health center in the villages. The study of population-averaged associations between malaria prevalence and covariates is useful for nationwide decision-making; for example, on supplying bed nets. The prediction of malaria prevalence in new locations is also of substantial interest. The previous literature suggests residual spatial dependence in these data (Diggle et al.,, 2002; Bai et al.,, 2014; Cattelan and Varin,, 2018).
@@
We choose the same set of variables as Cattelan and Varin, (2018), reproducing their results based on spatial generalized estimating equations, and comparing with bridge and Gaussian process random effect logistic models with an exponential correlation kernel. We use the same priors as in the simulation study (Gelman,, 2006; Gelman et al.,, 2008) except for the range parameter $\rho$. For the prior on the spatial range $\rho$ in the exponential correlation kernel $\mathcal{R}(s,s^{\prime})=\exp(\|s-s^{\prime}\|_{2}/\rho)$, we assign a uniform prior $\rho\sim\mathrm{Unif}(0.01,100)$ for the bridge and Gaussian process random effects models. We run three Markov chains with 11,000 iterations, where the first 1,000 samples are discarded as burn-in and 10,000 samples are saved without thinning. The running time based on a pure R implementation is about 2 mins for the bridge process random effect model and about 1.5 mins for the Gaussian process random effect model under the Apple M1 3.20GHz CPU environment.
@@
The results are summarized in Table 3. Trace plots of the bridge process and Gaussian process random effect models are shown in Figure A.4, both showing excellent convergence. The site-specific estimates using the bridge and Gaussian process random effect models give almost identical results. From the bridge process model, the posterior mean of the attenuation parameter is $\hat{\phi}=0.75$ with 95% credible interval $(0.46,0.97)$. The spatial generalized estimating equation and the proposed method give similar results for population-averaged estimates, except for the intercept and greenness. Specifically, the significance of the intercept and greenness coefficients at the 0.05 level differs across two methods. This discrepancy can be partly explained by simulation results, where the coverage of the intercept estimate $\hat{\beta}^{M}_{0}$ under the true $\beta_{0}^{M}=0$ is lower than the nominal level with the spatial generalized estimating equation method but not with the proposed method.
In addition to the estimated population-averaged and site-specific coefficients, we have posterior mean $\hat{\rho}=33.3(14.2,47.2)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the bridge process, and posterior mean $\hat{\rho}=29.9(13.4,40.1)$ and posterior mode $\hat{\rho}^{\mathrm{mode}}=13.5$ for the Gaussian process random effect models, where parentheses correspond to the posterior interquartile range. The widely applicable information criterion (Gelman et al.,, 2014), conditional on random effects, gives 2329.5 (standard error 40.2) for the bridge process random effect model and 2326.3 (standard error 40.0) for the Gaussian process random effect model. Although the Gaussian process model shows slightly better predictive performance, the difference is negligible when considering the benefit of dual interpretation offered by the bridge process model.
6 Discussion
Although we have focused on spatial settings, the proposed bridge process can also be applicable to modeling longitudinal binary data with time-correlated random effects (Parzen et al.,, 2011). For example, for discrete-time indices $t$ and $t^{\prime}$, one can employ an AR(1)-type correlation kernel $\mathcal{R}(t,t^{\prime})=\varrho^{|t-t^{\prime}|}$ for some $\varrho\in(-1,1)$. For such a case, the corresponding posterior inference algorithm can be appropriately adjusted to accommodate multiple realizations of bridge processes with a common parameter $(\phi,\varrho)$. Furthermore, we anticipate that recent advances in fast approximate Bayesian methods for non-Gaussian latent models (Cabral et al.,, 2024) could also be integrated with the bridge process, leveraging the mixture representation in Theorem 1.
@@ Acknowledgement
This research was partially supported by the National Institutes of Health (grant ID R01ES035625), by the European Research Council under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 856506), by the National Science Foundation (NSF IIS-2426762), and by the Office of Naval Research (N00014-21-1-2510). The authors thank Amy Herring and Georgia Papadogeorgou for helpful discussions.
Software
Code to reproduce the analyses is available at https://github.com/changwoo-lee/spbridge.
@@
Section A.1 contains proofs of the statements presented in the main article. Section A.2 describes the posterior inference algorithm and associated computational strategies. Section A.3 contains details of the multivariate bridge distribution, including a mixture representation. Section A.4 includes an additional figure corresponding to the analysis of the Gambia malaria data.
@@ A.2.2 Step 2
In all data analysis examples where the spatial range parameter $\rho$ is sampled together with $\phi$, we set $L=20$ and use coordinatewise logit transform to map parameters $(\phi,\rho)\in(0,1)\times(0.001,0.3)$ to $\mathbb{R}^{2}$, and utilize Metropolis-Hastings proposal adaptation settings as in Haario et al., (2001) to determine proposal distribution $q(\phi^{\star},\rho^{\star}\mid\phi,\rho)$.
@@ A.2.4 Details of scalable computation with low-rank dependence structure
We describe in detail how Algorithm 1 becomes scalable with the correlation kernel (8) with a low-rank structure. Denoting $R_{nq}=[\mathcal{R}(s_{i},\tilde{s}_{k})]_{i=1,k=1}^{n,q}$ as an $n\times q$ matrix and $R_{qn}=R_{nq}^{\mathrm{\scriptscriptstyle T}}$, we have a low-rank structured correlation matrix for $n$ realizations $\tilde{R}=[\tilde{R}(s_{i},s_{i^{\prime}})]_{i,i^{\prime}=1}^{n,n}=R_{nq}R_{qq}^{-1}R_{qn}+D_{nn}$ where $D_{nn}$ is a diagonal matrix with elements $1-r(s_{i})^{\mathrm{\scriptscriptstyle T}}R_{qq}^{-1}r(s_{i})$ for $i=1,\dots,n$.
First, step 1 of Algorithm 1 involves $n\times n$ matrix inversion in equation (A.1). This can be reduced to a $q\times q$ matrix inversion problem using the Woodbury matrix identity,
@@ A.3 Details of multivariate bridge distribution
A.4 Additional figure for Gambia malaria study